A 4000 investment earns interest at 4%/ yr compunded quarterly. How long will it take to double the value of the investment. Please show me 4 steps
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

8000=4000(1+(.04/4))^4t, where t is number of years
2=(1+0.01)^4t
ln both sides
ln 2=4t ln 1.01 rule of log
t=ln2/(4 ln(1.01)
=17.42 years
Rule of 70
70/rate in percent = time in years to double.
Here, it would yield 70/4=17.5 years.

Answer by Alan3354(69443)

4%/ yr compunded quarterly. How long will it take to double the value of the investment.
-------------
The amount doesn't matter.
4%/year ---> 1%/quarter
===============
FV = future value
PV = present value
-----
FV = 2*PV = PV*(1.01)^q --- q = # of quarters
2 = 1.01^q
log(2) = log(1.01^q) = q*log(1.01)
q = log(2)/log(1.01)